OpenStudy (anonymous):

-3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

do you know how to do this @Mokeira ?

OpenStudy (anonymous):

@aajugdar help

OpenStudy (anonymous):

oh hey you have to create two new equations

OpenStudy (anonymous):

one by adding -3x + 7y = -16 and -9x + 5y = 16 and second by multiplying -3x + 7y = -16 and -9x + 5y = 16

OpenStudy (anonymous):

do i just add the x's and y's together

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

-3x + 7y +(-9x + 5y) = -16+16

OpenStudy (anonymous):

ok thank and how do i do part b

OpenStudy (anonymous):

in 2nd you have to solve original two equations as well as new equations and you have to show that values of x and y match in both systems of equations

OpenStudy (anonymous):

i don't know how to solve it

OpenStudy (anonymous):

@aajugdar

OpenStudy (anonymous):

oh? hmm okay i will help

OpenStudy (anonymous):

-3x + 7y = -16 -9x + 5y = 16 so you have two equations here there are many ways to solve this but i will tell you elimination method okay?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

in elimination method what you have to do is either eliminate x compnent or y component so in the equations -3x + 7y = -16 -9x + 5y = 16 suppose we want to eliminate x component now we have -3x and -9x as x components so we multiply first equation by -3 in order to eliminate those so you get -3(-3x+7y=-16) 9x-21y=48

OpenStudy (anonymous):

so now you have two equations 9x-21y=48 and -9x + 5y = 16

OpenStudy (anonymous):

now add these two equations 9x and -9x will be canceled -21y+5y = -16y 48+16=64 so we get -16y = 64 so y= 64/-16 y= -4

OpenStudy (anonymous):

now consider this equation -9x + 5y = 16 we know that value of y is -4 put this value in equation -9x+5(-4) =16 -9x-20=16 -9x=36 x=-4 so you have values of x and y by elimination method any doubts?

OpenStudy (anonymous):

thanks a lot helped me very much!!!!!!!!!

OpenStudy (anonymous):

you are welcome :)